When a block of mass 0.5 kg is suspended from the end of a spring, it stretches by 0.02 m. (a) Find the force constant if the spring. (b) The block is given by a sharp impulse, so that it moves upwards with a velocity of ✓20 ms-1. How high will it rise?

Grade:10

1 Answers

Spartan 007
6 years ago
A-) F=-kx mg=-k(0.02) -k=0.5*10/0.02 k=-250N/m. 77777777777777777777777777777777777777777777777777777777777777777777